Combine the terms if possible or enter DNE if they cannot be combined: - 4340 – 67320
Answer #1

Solution:-

Given that,

-4\sqrt[3]{40}-6\sqrt[3]{320}

We know that,

Prime\; \; factorization\; \; of\; \; 40:\; 2^3\times 5\; \; \; \; (i.e\; \; 2\times 2\times 2\times 5=40)

Prime\; \; factorization\; \; of\; \; 320:\; 2^6\times 5\; \; \; \; (i.e\; \; 2\times 2\times 2\times 2\times 2\times 2\times 5=320)

Therefore, the above expression can be rewritten as -

-4\sqrt[3]{40}-6\sqrt[3]{320}

-4\sqrt[3]{2^3\times 5}-6\sqrt[3]{2^6\times 5}

Applying radical rule:

  • \sqrt[n]{ab}= \sqrt[n]{a}\times \sqrt[n]{b}

Hence,

-4\sqrt[3]{2^3\times 5}-6\sqrt[3]{2^6\times 5}

-4[\sqrt[3]{2^3}\times \sqrt[3]{5}]-6[\sqrt[3]{2^6}\times \sqrt[3]{5}]

Applying radical rule:

  • \sqrt[n]{a^n} = a
  • \sqrt[n]{a^m} = a^{\frac{m}{n}}

Therefore,

-4[\sqrt[3]{2^3}\times \sqrt[3]{5}]-6[\sqrt[3]{2^6}\times \sqrt[3]{5}]

-4[2\times \sqrt[3]{5}]-6[2^{\frac{6}{3}}\times \sqrt[3]{5}]

-4[2\times \sqrt[3]{5}]-6[2^2\times \sqrt[3]{5}]

-4[2\times \sqrt[3]{5}]-6[4\times \sqrt[3]{5}]

-8\sqrt[3]{5}-24\sqrt[3]{5}

Adding similar elements -

-8\sqrt[3]{5}-24\sqrt[3]{5}

\boldsymbol{-32\sqrt[3]{5}}

Hence,

-4\sqrt[3]{40}-6\sqrt[3]{320} = \boldsymbol{-32\sqrt[3]{5}}